OpenStudy (anonymous):

Can someone convert 2 yd 1ft to inches and show me the steps

OpenStudy (anonymous):

There are 3 feet in a yard, so two yards is 3*2=6 feet, so the length you are talking about is 6+1 feet or 7 feet. There are 12 inches in a foot, so we have 12*7=84 inche. For more solutions to textbook problems, be sure to check out http://www.slader.com/s/eWFrZXlnbGVl . I submit a lot of content there as well.

OpenStudy (anonymous):

since in 3 foot =1 yard and in 1 foot there is 12 inches you use dimensional analysis you know you have 2 yrd and 1 foot 2rd+(1/3)yrd=(7/3)yrd 7/3yrd(3ft/1yrd)(12inches/1ft) yrds cancel and feet cancel and your left with inches

OpenStudy (anonymous):

1 yard = 3 feet so 2 yard = 3*2 =6 feet 1 feet = 12 inch so 6 feet = 12*6 =72 inches